OpenStudy (anonymous):

The block B of mass m is on the smooth inclined surface of wedge A and mass M on a smooth floor what should be the magnitude of the horizontal force F to be applied on A so that B will not slide on the inclined surface?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ F= m g \tan(\theta) \] Am I right?

OpenStudy (anonymous):

 \[ Nsin\theta = ma = m \cdot \frac{F}{m + M} \] \[ Ncos\theta = mg \] Divide both the equations: \[ tan\theta = \frac{F}{g(m + M)} \] \[ F = (M + m)g \cdot tan\theta \] Should be correct (though not sure)